After reading this chapter, it isnt surprising that youre becoming an investment wizard. With your newfound expertise, you
Answer #1

1)

Total return = (ending price - Purchasing price + dividends) / purchasing price

= (57.41 - 44.51 + 0.51) / 44.51

= 30.13%

2)

tax on dividend = 15%

dividend = 0.51

after tax dividend = 0.51*(1-15%) = 0.4335

Return = (ending price - Purchasing price + after tax dividends) / purchasing price

= (57.41 - 44.51 + 0.4335) / 44.51

= 29.96%

3)

capital gain = 57.41 - 44.51 = 12.9

Tax rate = 15%

after tax capital gain = 12.9*(1 - 0.15) = 10.965

Return = (after tax capital gain + after tax dividends) / purchasing price

= (10.965 + 0.4335) / 44.51

= 25.61%